# Uniform Norm Error Estimates for 2D Turning Point Problem

### v1

@@ 3 Estimates for Discrete Green’s function
$\leq C\sqrt{\varepsilon}\left(\frac{\varepsilon\log(1/\varepsilon)}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\;N$
$\leq C\frac{\varepsilon^{3}(\log(1/\varepsilon))^{5/2}}{N}\leq C\frac{(\log(1/\varepsilon))^{5/2}}{N}.$ (3.5)
Now, consider the third term of the R.H.S. of (3.1) in the $x$-direction $\varepsilon((R-IR)_{x},G_{mn,x})_{\Omega_{f,y}}$. We obtain,
$\varepsilon((R-IR)_{x},G_{mn,x})_{\Omega_{f,y}}$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{R}_{xyy}\|_{\Omega_{f,y}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{f,y}}}$
@@
$\leq C\sqrt{\varepsilon}\left(\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{N}\right)^{2}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;N$
$\leq C\frac{\varepsilon^{7/4}(\log(1/\varepsilon^{3/2}))^{5/2}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{5/2}}{N}.$ (3.6)
Now, consider the third term of the R.H.S. of (3.1) in the $y$-direction $\varepsilon((R-IR)_{y},G_{mn,y})_{\Omega_{f,y}}$. We obtain,
$\varepsilon((R-IR)_{y},G_{mn,y})_{\Omega_{f,y}}$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{f,y}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{f,y}}}$
@@
$\leq C\sqrt{\varepsilon}\left(\frac{1}{N}\right)^{2}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;N$
$\leq C\frac{\varepsilon^{3/4}(\log(1/\varepsilon^{3/2}))^{1/2}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{1/2}}{N}.$ (3.7)
@@
Now, consider the fourth term of the R.H.S. of (3.1) $\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega_{c}}$. We obtain,
$\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega_{c}}$ $\leq\varepsilon^{1/2}\|\nabla(R-IR)\|\varepsilon^{1/2}\|G_{mn,x}\|_{\Omega_{c}}$
$\leq\varepsilon^{1/2}CN^{-2}N$
$\leq C\varepsilon^{1/2}\frac{1}{N}\leq\frac{C}{N}.$ (3.8)
@@
Applying similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then we have
$\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega}\leq C\frac{(\log(1/\varepsilon^{3/2})^{3}}{N}.$
@@
$\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega}\leq\frac{C(\log(1/\varepsilon^{3/2}))^{3}}{N}$
Lemma 4.2 When $\alpha\geq 3/2$ and the source point $(x_{m},y_{n})$ lies in $\Omega_{c}\cup\Omega_{f,x}$, then $\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})_{\Omega}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}$.
@@
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{c}$.
$\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})$ $=\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})_{\Omega_{f,xy}}+\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})_{\Omega_{f,x}}$
$+\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})_{\Omega_{f,y}}+\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})_{\Omega_{c}}$ (3.9)

### v2

@@ 4 Estimates related to diffusion terms
$\leq C\sqrt{\varepsilon}\left(\frac{\varepsilon\log(1/\varepsilon)}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\frac{\varepsilon^{3}(\log(1/\varepsilon))^{3}}{N^{3/2}}\leq C\frac{(\log(1/\varepsilon))^{3}}{N^{3/2}}.$ (4.5)
Now, estimate the third term of the R.H.S. of (4.1) in the $x$-direction is $|\varepsilon((R-IR)_{x},G_{mn,x})_{\Omega_{f,y}}|$. By using Theorem 1.2, Theorem 4.1 and furthermore, we obtain
$|\varepsilon((R-IR)_{x},G_{mn,x})_{\Omega_{f,y}}|$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{R}_{xyy}\|_{\Omega_{f,y}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{f,y}}}$
@@
$\leq C\sqrt{\varepsilon}\left(\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{N}\right)^{2}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\frac{\varepsilon^{7/4}(\log(1/\varepsilon^{3/2}))^{3}}{N^{3/2}}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N^{3/2}}.$ (4.6)
Now, estimate the third term of the R.H.S. of (4.1) in the $y$-direction is $|\varepsilon((R-IR)_{y},G_{mn,y})_{\Omega_{f,y}}|$. By using Theorem 1.2, Theorem 4.1 and furthermore, we obtain
$|\varepsilon((R-IR)_{y},G_{mn,y})_{\Omega_{f,y}}|$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{f,y}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{f,y}}}$
@@
$\leq C\sqrt{\varepsilon}\left(\frac{1}{N}\right)^{2}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\frac{\varepsilon^{3/4}\log(1/\varepsilon^{3/2})}{N^{3/2}}\leq C\frac{\log(1/\varepsilon^{3/2})}{N^{3/2}}.$ (4.7)
@@
Now, estimate the fourth term of the R.H.S. of (4.1) $|\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega_{c}}|$. By using Theorem 1.2, Theorem 4.1 and furthermore, we obtain
$|\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega_{c}}|$ $\leq\varepsilon^{1/2}\|\nabla(R-IR)\|\varepsilon^{1/2}\|G_{mn,x}\|_{\Omega_{c}}$
$\leq\varepsilon^{1/2}CN^{-2}\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\varepsilon^{1/2}\frac{(\log(1/\varepsilon))^{1/2}}{N^{3/2}}\leq\frac{C(\log(1/\varepsilon))^{1/2}}{N^{3/2}}.$ (4.8)
@@
Applying similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{c}$, then we have
$|\varepsilon(\nabla(R-IR),\nabla G_{mn})|\leq C\frac{(\log(1/\varepsilon^{3/2})^{3}}{N}.$
Hence, for the source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, we have
$|\varepsilon(\nabla(R-IR),\nabla G_{mn})|\leq\frac{C(\log(1/\varepsilon^{3/2}))^{3}}{N}$
Lemma 4.3
If the source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, then for $\alpha\geq 3/2$ following estimate hold
$|\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})|\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}.$
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{f,x}$ and decompose the domain in four parts, we have
$|\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})|$ $\leq|\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})_{\Omega_{f,xy}}|+|\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})_{\Omega_{f,x}}|$
$+|\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})_{\Omega_{f,y}}|+|\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})_{\Omega_{c}}|$ (4.9)
Now, estimate the first term of the R.H.S. of (4.9) in the $x$-direction is $|\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{f,xy}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.3.1 and furthermore, we obtain
$|\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{f,xy}}|$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{F_{1}}_{xyy}\|_{\Omega_{f,xy}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{f,xy}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{\log(1/\varepsilon^{3/2})}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\varepsilon^{5/4}\frac{(\log(1/\varepsilon^{3/2}))^{7/2}}{N^{3/2}}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{7/2}}{N^{3/2}}.$ (4.10)
Now, estimate the first term of the R.H.S. of (4.9) in the $y$-direction is $|\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,xy}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.3.1 and furthermore, we obtain
$|\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,xy}}|$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{f,xy}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{f,xy}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{\log(1/\varepsilon^{3/2})}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\varepsilon^{5/4}\frac{(\log(1/\varepsilon^{3/2}))^{7/2}}{N^{3/2}}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{7/2}}{N^{3/2}}.$ (4.11)
Now, estimate the second term of the R.H.S. of (4.9) in the $x$-direction $|\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{f,x}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.3 and furthermore, we obtain
$|\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{f,x}}|$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{F_{1}}_{xyy}\|_{\Omega_{f,x}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{f,x}}}$
$\leq C\sqrt{\varepsilon}\frac{1}{\varepsilon N^{2}}\sqrt{\varepsilon\log(1/\varepsilon)}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\frac{\log(1/\varepsilon)}{N^{3/2}}.$ (4.12)
Now, estimate the second term of the R.H.S. of (4.9) in the $y$-direction is $|\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,x}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.3.1 and furthermore, we obtain
$|\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,x}}|$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{f,x}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{f,x}}}$
$\leq C\sqrt{\varepsilon}\left(\frac{\log(1/\varepsilon)}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\varepsilon^{1/2}\frac{(\log(1/\varepsilon))^{3}}{N^{3/2}}\leq C\frac{(\log(1/\varepsilon))^{3}}{N^{3/2}}.$ (4.13)
Now, estimate the third term of the R.H.S. of (4.9) in the $x$-direction is $|\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{f,y}}|$. By using Theorem 1.2, Theorem 4.1, Lemma 3.3 and furthermore, we obtain
$|\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{f,y}}|$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{F_{1}}_{xyy}\|_{\Omega_{f,y}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{f,y}}}$
$\leq C\sqrt{\varepsilon}\frac{(\log(1/\varepsilon^{3/2}))^{2}}{N^{2}}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;\varepsilon^{\alpha}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\varepsilon^{\alpha+3/4}\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N^{3/2}}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N^{3/2}}.$ (4.14)
Now, estimate the third term of the R.H.S. of (4.9) in the $y$-direction is $|\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,y}}|$. By using Theorem 1.2, Theorem 4.1, Lemma 3.3 and furthermore, we obtain
$|\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{f,y}}|$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{f,y}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{f,y}}}$
$\leq C\sqrt{\varepsilon}\frac{1}{N^{2}}\sqrt{|\Omega_{f,y}|}\frac{1}{\varepsilon^{2}}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{f,y}})}\|G_{mn}\|_{1,\varepsilon}$
$\leq C\sqrt{\varepsilon}\frac{1}{N^{2}}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\frac{1}{\varepsilon^{2}}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{f,y}})}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\frac{\log(1/\varepsilon^{3/2})}{N^{3/2}}\varepsilon^{\alpha-3/2}\leq C\frac{\log(1/\varepsilon^{3/2})}{N^{3/2}}\quad\text{when}\;\alpha\geq 3/2.$ (4.15)
Now, estimate the fourth term of the R.H.S. of (4.9) in the $x$-direction is $|\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{c}}|$. By using Theorem 1.2, Theorem 4.1, Lemma 3.3 and furthermore, we obtain
$|\varepsilon((F_{1}-IF_{1})_{x},G_{mn,x})_{\Omega_{c}}|$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{F_{1}}_{xyy}\|_{\Omega_{c}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{c}}}$
$\leq C\frac{1}{N^{2}}\frac{1}{\sqrt{\varepsilon}}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{c}})}\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\varepsilon^{-1/2}\frac{(\log(1/\varepsilon))^{1/2}}{N^{3/2}}\varepsilon^{\alpha}\leq\frac{C(\log(1/\varepsilon))^{1/2}}{N^{3/2}}.\quad\text{when}\;\alpha\geq 1/2.$ (4.16)
Now, estimate the fourth term of the R.H.S. of (4.9) in the $y$-direction is $|\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{c}}|$. By using Theorem 1.2, Theorem 4.1, Lemma 3.3 and furthermore, we obtain
$|\varepsilon((F_{1}-IF_{1})_{y},G_{mn,y})_{\Omega_{c}}|$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{c}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{c}}}$
$\leq C\frac{1}{N^{2}}\frac{1}{\varepsilon^{3/2}}\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}({\Omega_{c}})}\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\frac{\varepsilon^{\alpha-3/2}}{N^{3/2}}(\log(1/\varepsilon))^{1/2}\leq\frac{C(\log(1/\varepsilon))^{1/2}}{N^{3/2}}.\quad\text{when}\;\alpha\geq 3/2.$ (4.17)
Applying similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{c}$, then we have
$|\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})|\leq C\frac{(\log(1/\varepsilon^{3/2})^{3}}{N}.$
@@
$|\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})|\leq C\frac{(\log(1/\varepsilon^{3/2})^{3}}{N}.$
Lemma 4.4
If the source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, then for $\beta\geq 1/3$ following estimate hold
$|\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})|\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}.$
@@
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{f,x}$ and decompose the domain in four parts, we have
$|\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})|$ $\leq|\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})_{\Omega_{f,xy}}|+|\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})_{\Omega_{f,x}}|$
$+|\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})_{\Omega_{f,y}}|+|\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})_{\Omega_{c}}|$ (4.18)